# Recent progresses in strange quark stars

### v1

@@ 5.1 GWs from binary strange star systems
Coalescence of binary systems is the most significant stellar GW sources. BH-BH, BH-NS, and NS-NS binary systems are common GW sources. Here we focus on binaries containing strange quark stars. According to the strange quark matter hypothesis (Witten, 1984; Farhi and Jaffe, 1984), pulsars may actually be strange stars. Therefore, we could have BH-SQS and SQS-SQS systems. The detection of the controversial GW190814 event, which involves a possible mass-gap compact object ( – ), has drawn a wide attention in the community. Although most researchers believe it could be a neutron star, the possibility that it is a strange star has also been discussed in numerous studies (Bombaci et al., 2021; Miao et al., 2021; Roupas et al., 2021; Lopes and Menezes, 2022; Oikonomou and Moustakidis, 2023). Interestingly, two frequently mentioned GW events, GW170817 and GW190425, are suggested to originate from SQS-SQS systems by Miao et al. (2021), Kumar et al. (2022), and Sagun et al. (2023). Furthermore, theoretical calculations of GW radiation from SQS-SQS binary systems have been preformed by Limousin et al. (2005), Gondek-Rosinska and Limousin (2008), and Bauswein et al. (2010). Moraes and Miranda (2014) even studied the possible existence of NS-SQS binaries and investigated their GW emission.
@@ 7 Concluding Remarks
Strange quark matter could be the true ground state of matter at extreme densities. Such a hypothesis need to be tested through astronomical investigations. Essentially, we should try to discriminate between neutron stars and strange stars through observations. In this article, we present a brief review on some recent progresses in this field. Various models describing quark confinement are introduced. The corresponding EOS derived from these models are presented and compared. By combining these EOSs with the TOV equations, we can calculate the inner structure of strange stars, deriving the mass-radius relation for the whole sequence of strange quark objects. Due to the different internal composition between neutron stars and strange stars, gravitational wave emissions from them are different, which could be potentially discerned through observations. Some special kinds of electromagnetic bursts could also be connected with strange stars and can be used as a probe of these exotic objects.
@@
X-LZ: conceptualization, investigation, validation, and writing–original draft. Y-FH: conceptualization, funding acquisition, validation, writing–review, and editing. Z-CZ: conceptualization, validation, writing–review, and editing.

### v2

@@ 6.1 GWs from binary strange star systems
Coalescence of binary systems is the most significant stellar GW sources. BH-BH, BH-NS, and NS-NS binary systems are common GW sources. Here we focus on binaries containing strange quark stars, such as BH-SQS and SQS-SQS systems. The detection of the controversial GW190814 event, which involves a possible mass-gap compact object ( – ), has drawn a wide attention in the community. Although most researchers believe it could be a neutron star, the possibility that it is a strange star has also been discussed in numerous studies (Bombaci et al., 2021; Miao et al., 2021; Roupas et al., 2021; Lopes and Menezes, 2022; Oikonomou and Moustakidis, 2023). Interestingly, two frequently mentioned GW events, GW170817 and GW190425, are suggested to originate from SQS-SQS systems by Miao et al. (2021), Kumar et al. (2022), and Sagun et al. (2023). Furthermore, theoretical calculations of GW radiation from SQS-SQS binary systems have been preformed by Limousin et al. (2005), Gondek-Rosinska and Limousin (2008), and Bauswein et al. (2010). Moraes and Miranda (2014) even studied the possible existence of NS-SQS binaries and investigated their GW emission.
@@ 8 Concluding Remarks
Strange quark matter could be the true ground state of matter at extreme densities. Such a hypothesis need to be tested through astronomical investigations. Essentially, we should try to discriminate between neutron stars and strange stars through observations. In this article, we present a brief review on some recent progresses in this field. Various models describing quark confinement are introduced. The corresponding EOS derived from these models are presented and compared. By combining these EOSs with the TOV equations, we can calculate the inner structure of strange stars, deriving the mass-radius relation for the whole sequence of strange quark objects. The tidal deformability and the Love number measured through gravitational wave observations may help diagnose the EOS. Radial and non-radial oscillations with different modes can also be used to probe the internal structure of strange stars. The properties of hybrid stars, in which quarks and hadrons may coexist, are also discussed. Some special kinds of electromagnetic bursts could also be connected with strange stars and can be used as a probe of these exotic objects.
@@
X-LZ: conceptualization, investigation, validation, and writing-original draft. Y-FH: conceptualization, funding acquisition, validation, writing-review, and editing. Z-CZ: conceptualization, validation, writing-review, and editing.